# ProLLM: Protein Chain-of-Thoughts Enhanced LLM for Protein-Protein Interaction Prediction

### v1

@@ 1 Introduction
Protein-protein interactions (PPIs) play an essential role in various biological processes of all living organisms, which are crucial for biomedical, genetic, and pharmaceutical research. Thus, numerous experimental methods have been proposed for PPI detection, such as yeast two-hybrid [1] and quantitative proteomics methods [2]. However, wet-lab methods for PPI prediction are often time-consuming and labor-intensive, highlighting the need for more precise and efficient computational tools.
In recent years, computational biology has developed rapidly. Methods such as the Convolutional Neural Network (CNN) and Graph Neural Network (GNN) have become powerful tools for studying protein interaction. CNN-based approaches like TAG-PPI [3] typically use pre-trained embedding models to convert the amino acid sequences in proteins into continuous vector representations, and then employ one-dimensional convolutional neural networks to extract features from the vectors for subsequent PPI tasks. Although CNN methods have shown some effectiveness in PPI prediction, they still have limitations due to their fixed receptive fields and the lack of well-defined spatial relationships in protein sequences, which limit the accuracy and interpretability of the predictions. GNN-based methods such as GNN-PPI [4] treat proteins as nodes and their relationships as edges, constructing a network composed of proteins, which better captures protein relationships and interactions, and outperforms CNNs in predicting protein interactions. However, while GNNs can effectively utilize network structural information, they are less effective at learning protein chains than transformer-based models  [5]. Furthermore, GNNs cannot fully capture and understand the relationships and dynamic changes in the signal passing process in real organisms, restricting their development for protein prediction  [6].
@@
Following the GNN and CNN methods, Large Language Models have also been applied to this PPI area, such as ProBert [7] and ProteinLM [8]. As long as these models can obtain a protein representation, we can use the direct cosine similarity of the representation or train an MLP to perform PPI prediction. However, these methods still cannot capture the chain relationships between proteins, such as the signaling pathways. Besides, previous literature only used LLMs as a feature extractor. Recently, using LLM as a link predictor has shown that it can better capture relational information between nodes in knowledge graph tasks and its performance surpasses traditional GNN baselines [9, 10, 11]. Therefore, it is promising to introduce LLM for protein-protein interaction (PPI) tasks, since the most important biological signal for PPI tasks is the chain relationships of proteins, i.e., the signaling pathways.
@@
The signaling pathway addresses the traditional method’s ignorance of global, non-physical connections between proteins. Signaling pathways typically start with an upstream protein that sends a biological signal through several intermediates to a downstream protein, hence requiring consideration of the cumulative effect of multiple protein interactions. This series of interactions form sequential chains. Therefore, we propose Protein Chain of Thought (ProCoT) to overcome the limitation in understanding signaling pathways and protein functions. ProCoT is a data format that simulates the signal transduction process using a thought-chain approach, thereby enabling the prediction of protein interactions in signaling pathway problems. CoT can express the thinking process step by step to form a reasoning chain [12], while ProCoT extends this principle further into the protein-related domain to simulate protein signaling pathways, giving LLMs a deeper insight into protein.
@@
Additionally, our approach addresses the issue of poor protein comprehension in LLMs by replacing the standard language model embedding with embedding infused with protein information. When we process the protein name in the prompt, we replace its original embedding by ProtTrans [13], because its embedding contains protein’s structual information. We also perform the instruction fine-tuning on the protein knowledge dataset to infuse protein domain knowledge into LLMs. Following these steps, the LLM acquires a robust ability to reason about the direct relationships between proteins, as demonstrated in Figure 1. It can provide answers to questions about protein relationships, which play a significant role in biological research.
@@
  • Experiments on four widely used PPI datasets (i.e., Human, SHS27K, SHS148K, and STRING) demonstrate that ProLLM outperforms graph-based methods such as GNN-PPI and SemiGNN-PPI. It also has better performance than LLM-based methods like InstructGLM. The micro-F1 scores on these 4 datasets are 91.05, 85.32, 87.66, 89.21, respectively.
@@ 2 Related Work
Protein-protein interactions (PPIs) are indispensable for all living organisms, and so many efforts have been made for PPI detection up to now. Yeast two-hybrid (Y2H) assays [14], synthetic lethality [15], quantitative proteomics [16], and mass spectrometry [17], are widely used for identifying PPIs. To be specific, Y2H assays explore binary PPIs in living cells, offering insight into protein functions and regulations, although labor-intensive and limited in genomic coverage. Synthetic lethality identifies essential gene pair interactions by revealing compensatory relationships when both are inactivated. Meanwhile, quantitative proteomics, especially through mass spectrometry, illuminates the dynamic nature of the interactome under various conditions. Although informative, these methods require significant labor, time, and resources.
@@
In the realm of machine learning, sequence-based approaches include Shen’s SVM method [18], which uses a 3-mer count vector from protein sequences as features and groups 20 amino acids into seven classes to handle synonymous mutations and reduce feature space dimensionality. SVM-based methods [19] and the ensemble model PCA-EELM (Principal Component Analysis-Ensemble Extreme Learning Machine) [20] utilize various types of protein sequence information for PPI prediction. In the domain of deep learning, DeepPPI [21] extracts a multitude of features from protein sequences and employs a dual deep neural network structure for feature fusion and prediction. Sun et al. [22] introduced a PPI predictor based on a stacked autoencoder, emphasizing the importance of sample balance. DPPI [23] and TAGPPI [3] further extend the application of convolutional neural networks and integrate text convolution networks with graph representation learning to enhance the accuracy of PPI predictions. GNNs have significantly advanced PPI predictions, improving our understanding of biological mechanisms. GNN-PPI [4] enhances inter-novel-protein prediction accuracy by utilizing protein relationships and a new evaluation approach. PT-GNN [24] integrates diverse data for link prediction, learning node features from sequence and structure. DeepRank-GNN [25] offers a modular, Python-packaged framework for GNN-based interaction pattern predictions. HIGH-PPI [26] introduces a hierarchical graph learning model for effective PPI prediction and molecular detail extrapolation. Geometric GNNs excel in modeling spatial intricacies, enhancing biomolecule prediction accuracy. Geo-PPI [27] utilizes self-supervised learning for geometric protein structure representations, excelling in detailing protein interactions. mmCSM-PPI [28] captures multifaceted features for mutation impact predictions on protein interactions. MAPE-PPI [29] defines the microenvironment of amino acid residues in proteins and encodes it into discrete codes, which can capture the connections between different microenvironments, enhancing the prediction accuracy of PPI.
@@
Recent advances in large language models, such as BERT [30], GPT [31], LLaMA [32], and T5 [33], have significantly advanced the field of Natural Language Processing (NLP) to new heights. These models, having been trained on extensive textual corpora, exhibit exceptional capabilities across a diverse range of NLP applications. Inspired by LLMs, Protein Large Language Models(PLMs) pre-trained on large-scale protein sequences have emerged, such as ESM [34], ProtTrans [13] and ProteinBert [13]. PLMs provide a better representation of protein sequences by converting the protein sequences into the form of high-dimensional numerical vectors, known as embedding. With the protein sequences captured by the PLMs, the performances on diverse downstream tasks, such as structure prediction [35], subcellular localization prediction [36], single peptide prediction [37] and N-linked glycosylation sites prediction [38], have been transformed. It can be expected that PLMs will assist in PPI prediction tasks. ProtLLM [10] utilizes a dynamic protein mounting mechanism, a protein-as-word language modeling approach, and the InterPT dataset for pre-training, enabling it to handle complex inputs and achieve superior performance on various proteins-related tasks. However, ProtLLM is used for general protein tasks, and it is not used for PPI tasks exactly. The methodology of training these LLMs to convert text inputs to desired text outputs positions them as particularly advantageous for tasks such as generative link prediction [9, 11]. In such tasks, the model is tasked with inferring and generating the relationship between two entities based on provided textual cues. Moreover, the extensive pre-training phase enables LLMs to exhibit a remarkable capacity for generalization. This capacity allows them to effectively tackle and respond to tasks or prompts that were not explicitly covered during their initial training [39]. In addition to the outlined capabilities, the inherent flexibility and generalization potential of LLMs suggest that their applicability extends well beyond the conventional boundaries of NLP tasks [40]. Specifically, their proficiency in generalizing from expansive pre-training sessions paves the way for their application in fields like bioinformatics and complex network analysis.
@@ 3 Proposed ProLLM Framework
In this section, we will introduce the implementation details of ProLLM, a framework designed to transform protein interaction data into ProCoT format natural language descriptions in order to simulate protein signaling pathways. By translating the structure and relationships of proteins into natural language, we effectively transform protein interaction problems into natural language processing (NLP) tasks. To enhance the understanding of proteins by ProLLM, we directly integrate the protein vectors generated from ProtTrans [13] into our model to replace the original word embedding in the protain name’s place. This approach allows our model to understand and utilize the biological attributes of proteins more accurately when predicting protein interactions. The ProLLM process is shown in Figure 3.
@@ 3.1.1 Protein Data to Natural Language in CoT Format
(Protein Signaling Pathway) In cellular signal transduction, receptor protein $\mathcal{R}\text{c}$ is the beginning of the signal pathway, and it can interact with downstream signaling proteins $\mathcal{S}\text{p}$. $\mathcal{S}\text{p}$ is responsible for continuing to transmit signals to effector protein $\mathcal{E}\text{p}$ inside the cell. After the $\mathcal{E}\text{p}$ response the signal, it will trigger specific biological effects and activate secondary messengers $\mathcal{M}$. $\mathcal{M}$ may be the end of this signal level, or it may open second signal hierarchy. If in the second signal level, the $\mathcal{M}$ will replace the $\mathcal{R}\text{c}$ in the initial signal hierarchy as the starting point. The signal will continue to propagate from $\mathcal{R}\text{c}$ through new signaling proteins, effector proteins, and then generate new secondary messengers. After this stage, the propagation of the signal repeats the structure of the second signal level until the signal is interrupted.
We aim to enhance the model’s understanding of biological signaling pathways, enhancing it’s ability to learn and reason interactions within complex signal networks. We create prompts in the ProCoT (Protein Chain of Thought) format to incrementally decompose the signal transduction process like Definition 2, simulating real pathways of signal propagation. Specifically, we clarify the structure knowledge within the signaling pathway, such as signaling proteins and effector proteins. We then design rules for signal transmission across different levels to simulate the iterative process of signal transduction in proteins. Finally, we also design signal interruptions to simulate the continuity of protein transmission in real life.
We use hard code to convert protein interaction information as Definition 1 in dataset into ProCoT format natural language to imitate protein signaling pathways as Definition 2. Figure 6 in the appendix is an example of ProCoT. The answer to this is "The relationship is activation".
@@ 3.1.2 Training on ProCot Format Dataset
In biology, co-expression and co-localization refer to the phenomenon in which proteins that are often expressed or located together in the cell tend to participate in the same or interconnected biological processes [41]. Thus, biologists frequently use known protein information to infer unknown protein interactions. Based on the principles of co-expression and co-localization in biology, we have formulated our training strategy. After constructing our ProCoT training data using DFS, we will obtain a circular protein interaction as in Figure 4. We mask the relationship between the initial and final proteins in the signaling pathway and let our model predict this relationship.
@@
3.1.3 Why ProCot Works: Biological Intuition behind ProCot:
In this section, we will explore ProCoT through the lenses of biology and AI. (1) Simulating Biological Signaling Pathways: The biological signaling pathway is a series of ordered, rule-based interaction processes. During the process, the output of each step serves as the input for the next step, forming a highly organized information transmission chain. This pattern is consistent with the methods of handling serialized information in natural language processing (NLP). Additionally, Flan-T5’s self-attention blocks can capture the indirect relationships between proteins that are distant and do not have direct interactions. It is crucial for understanding biological signal transduction processes that involve multiple steps and complex intermediate links. (2) Signal Interruption Mechanism: This mechanism mimics protein adaptability. Because the signaling between proteins needs to be constantly interrupted to ensure the cell’s accurate response to external changes. This mechanism we designed can satisfy the complex feedback mechanisms and regulatory networks within the cell. Overall, our ProCoT design follows biological principles.
@@ 3.2 Enhancing Protein Sequence and Function Comprehension
In our methodology, we enhance the model’s understanding of protein sequence by replacing the T5 embedding with the ProTrans embedding vectors, and also perform instruction fine-tuning to enable the model to learn protein functions.

### v2

@@ 1 Introduction
Protein-protein interactions (PPIs) play an essential role in various biological processes of all living organisms, which are crucial for biomedical, genetic, and pharmaceutical research. Thus, numerous experimental methods have been proposed for PPI detection, such as yeast two-hybrid (Ito et al., 2001) and quantitative proteomics methods (Rotilio et al., 2012). However, wet-lab methods for PPI prediction are often time-consuming and labor-intensive, highlighting the need for more precise and efficient computational tools.
In recent years, computational biology has developed rapidly. Methods such as the Convolutional Neural Network (CNN) and Graph Neural Network (GNN) have become powerful tools for studying protein interaction. CNN-based approaches like TAG-PPI (Song et al., 2022) typically use pre-trained embedding models to convert protein sequences into numerical vector representations, and then employ one dimensional convolutional neural networks to extract features from the vectors for subsequent PPI tasks.
Although CNN methods have shown some effectiveness in PPI prediction, they still have limitations due to their fixed receptive fields and the lack of well-defined spatial relationships in protein sequences, which limit the accuracy and interpretability of the predictions. GNN-based methods such as GNN-PPI (Lv et al., 2021) treat proteins as nodes and their relationships as edges, constructing a network composed of proteins, which better captures protein relationships and interactions, and outperforms CNNs in predicting protein interactions. However, while GNNs can effectively extract network structural information, they neglected the non-physical connections between two proteins without direct physical interactions, resulting in the worse performance in learning protein chains than transformer-based models  (Zhou et al., 2020). Furthermore, GNNs cannot fully capture the relationships and dynamic changes in the signal passing process of living organisms, restricting their performance for PPI prediction  (Zhou et al., 2023).
@@
Following the GNN and CNN methods, Large Language Models have also been applied to this PPI area, such as ProBert (Elnaggar et al., 2021a) and ProteinLM (Xiao et al., 2021). As long as these models can obtain a protein representation, we can use the direct cosine similarity of the representation or train an MLP to perform PPI prediction. However, these methods still cannot capture the chain relationships between proteins, such as the signaling pathways. Besides, previous literature only used LLMs as a feature extractor. Recently, using LLM as a link predictor has shown that it can better capture relational information between nodes in knowledge graph tasks and its performance surpasses traditional GNN baselines (Ye et al., 2024; Zhuo et al., 2024; Shu et al., 2024). Therefore, it is promising to introduce LLM for protein-protein interaction (PPI) tasks, since the most important biological signal for PPI tasks is the chain relationships of proteins, i.e., the signaling pathways.
@@
The signaling pathway addresses the traditional method’s ignorance of global, non-physical connections between proteins. Signaling pathways typically start with an upstream protein that sends a biological signal through several intermediates to a downstream protein, hence requiring consideration of the cumulative effect of multiple protein interactions. This series of interactions form sequential chains. Therefore, we propose Protein Chain of Thought (ProCoT) to overcome the limitation in understanding signaling pathways and protein functions. ProCoT is a data format that simulates the signal transduction process using a thought-chain approach, thereby enabling the prediction of protein interactions in signaling pathway problems. CoT can express the thinking process step by step to form a reasoning chain (Jin et al., 2024b), while ProCoT extends this principle further into the protein-related domain to simulate protein signaling pathways, giving LLMs a deeper insight into protein.
@@
Additionally, our approach addresses the issue of poor protein comprehension in LLMs by replacing the standard language model embedding with embedding infused with protein information. When we process the protein name in the prompt, we replace its original embedding by ProtTrans (Elnaggar et al., 2021b), because its embedding contains protein’s structual information. We also perform the instruction fine-tuning on the protein knowledge dataset to infuse protein domain knowledge into LLMs. Following these steps, the LLM acquires a robust ability to reason about the direct relationships between proteins, as demonstrated in Figure 1. It can provide answers to questions about protein relationships, which play a significant role in biological research.
@@
  • Experiments on four widely used PPI datasets (i.e., Human, SHS27K, SHS148K, and STRING) demonstrate that ProLLM outperforms graph-based methods such as GNN-PPI and SemiGNN-PPI. It also has better performance than LLM-based methods like InstructGLM. The micro-F1 scores on these 4 datasets are 91.05, 78.09, 87.66, 89.21, respectively.
@@ 2 Related Work
Protein-protein interactions (PPIs) are indispensable for all living organisms, and so many efforts have been made for PPI detection up to now. Yeast two-hybrid (Y2H) assays (Brückner et al., 2009), synthetic lethality (O’Neil et al., 2017), quantitative proteomics (Wilm, 2009), and mass spectrometry (Mann et al., 2001), are widely used for identifying PPIs. To be specific, Y2H assays explore binary PPIs in living cells, offering insight into protein functions and regulations, although labor-intensive and limited in genomic coverage. Synthetic lethality identifies essential gene pair interactions by revealing compensatory relationships when both are inactivated. Meanwhile, quantitative proteomics, especially through mass spectrometry, illuminates the dynamic nature of the interactome under various conditions. Although informative, these methods require significant labor, time, and resources.
@@
In the realm of machine learning, sequence-based approaches include Shen’s SVM method (Shen et al., 2007), which uses a 3-mer count vector from protein sequences as features and groups 20 amino acids into seven classes to handle synonymous mutations and reduce feature space dimensionality. SVM-based methods (Guo et al., 2008) and the ensemble model PCA-EELM (Principal Component Analysis-Ensemble Extreme Learning Machine) (You et al., 2013) utilize various types of protein sequence information for PPI prediction. In the domain of deep learning, DeepPPI (Du et al., 2017) extracts a multitude of features from protein sequences and employs a dual deep neural network structure for feature fusion and prediction. Sun et al. (Sun et al., 2017) introduced a PPI predictor based on a stacked autoencoder, emphasizing the importance of sample balance. DPPI (Hashemifar et al., 2018) and TAGPPI (Song et al., 2022) further extend the application of convolutional neural networks and integrate text convolution networks with graph representation learning to enhance the accuracy of PPI predictions. GNNs have significantly advanced PPI predictions, improving our understanding of biological mechanisms. GNN-PPI (Lv et al., 2021) enhances inter-novel-protein prediction accuracy by utilizing protein relationships and a new evaluation approach. PT-GNN (Long et al., 2022) integrates diverse data for link prediction, learning node features from sequence and structure. DeepRank-GNN (Réau et al., 2023) offers a modular, Python-packaged framework for GNN-based interaction pattern predictions. HIGH-PPI (Gao et al., 2023) introduces a hierarchical graph learning model for effective PPI prediction and molecular detail extrapolation. Geometric GNNs excel in modeling spatial intricacies, enhancing biomolecule prediction accuracy. Geo-PPI (Liu et al., 2021) utilizes self-supervised learning for geometric protein structure representations, excelling in detailing protein interactions. mmCSM-PPI (Rodrigues et al., 2021) captures multifaceted features for mutation impact predictions on protein interactions. MAPE-PPI (Wu et al., 2024) defines the microenvironment of amino acid residues in proteins and encodes it into discrete codes, which can capture the connections between different microenvironments, enhancing the prediction accuracy of PPI.
@@
Recent advances in large language models, such as BERT (Devlin et al., 2018), GPT (Peng et al., 2023), LLaMA (Touvron et al., 2023), and T5 (Raffel et al., 2020), have significantly advanced the field of Natural Language Processing (NLP) to new heights. These models, having been trained on extensive textual corpora, exhibit exceptional capabilities across a diverse range of NLP applications (Shengyuan et al., 2024; Jin et al., 2024a; Fan et al., 2024; Hua et al., 2024). Inspired by LLMs, Protein Large Language Models(PLMs) pre-trained on large-scale protein sequences have emerged, such as ESM (Hsu et al., 2022), ProtTrans (Elnaggar et al., 2021b) and ProteinBert (Elnaggar et al., 2021b). PLMs provide a better representation of protein sequences by converting the protein sequences into the form of high-dimensional numerical vectors, known as embedding. With the protein sequences captured by the PLMs, the performances on diverse downstream tasks, such as structure prediction (Lin et al., 2023), subcellular localization prediction (Thumuluri et al., 2022), single peptide prediction (Teufel et al., 2022) and N-linked glycosylation sites prediction (Hou et al., 2023), have been transformed. It can be expected that PLMs will assist in PPI prediction tasks. ProtLLM (Zhuo et al., 2024) utilizes a dynamic protein mounting mechanism, a protein-as-word language modeling approach, and the InterPT dataset for pre-training, enabling it to handle complex inputs and achieve superior performance on various proteins-related tasks. However, ProtLLM is used for general protein tasks, and it is not used for PPI tasks exactly. The methodology of training these LLMs to convert text inputs to desired text outputs positions them as particularly advantageous for tasks such as generative link prediction (Ye et al., 2024; Shu et al., 2024). In such tasks, the model is tasked with inferring and generating the relationship between two entities based on provided textual cues. Moreover, the extensive pre-training phase enables LLMs to exhibit a remarkable capacity for generalization. This capacity allows them to effectively tackle and respond to tasks or prompts that were not explicitly covered during their initial training (Wei et al., 2022). In addition to the outlined capabilities, the inherent flexibility and generalization potential of LLMs suggests that their applicability extends well beyond the conventional boundaries of NLP tasks (Yang et al., 2023). Specifically, their proficiency in generalizing from expansive pre-training sessions paves the way for their application in fields like bioinformatics and complex network analysis.
@@ 3 Proposed ProLLM Framework
In this section, we will introduce the implementation details of ProLLM, a framework designed to transform protein interaction data into ProCoT format natural language descriptions to simulate protein signaling pathways. By translating the structure relationships of proteins into natural language, we effectively transform protein interaction problems into natural language processing (NLP) tasks. To enhance the understanding of proteins by ProLLM, we directly integrate the protein vectors generated from ProtTrans (Elnaggar et al., 2021b) into our model to replace the original word embedding in the protein name’s place. This approach allows our model to understand and utilize the biological attributes of proteins when predicting protein interactions. The ProLLM process is shown in Figure 3.
@@ 3.1.1 Protein Data to Natural Language in CoT Format
(Protein Signaling Pathway) In cellular signal transduction, receptor protein $\mathcal{R}\text{c}$ is the beginning of the signal pathway, and it can interact with downstream signaling proteins $\mathcal{S}\text{p}$. $\mathcal{S}\text{p}$ is responsible for continuing to transmit signals to effector protein $\mathcal{E}\text{p}$ inside the cell. After the $\mathcal{E}\text{p}$ response to the signal, it will trigger specific biological effects and activate secondary messengers $\mathcal{M}$. $\mathcal{M}$ may be the end of this signal level, or it may open a second signal hierarchy. If in the second signal level, the $\mathcal{M}$ will replace the $\mathcal{R}\text{c}$ in the initial signal hierarchy as the starting point. The signal will continue to propagate from $\mathcal{R}\text{c}$ through new signaling proteins, and effector proteins, and then generate new secondary messengers. After this stage, the propagation of the signal repeats the structure of the second signal level until the signal is interrupted.
We aim to improve the model’s understanding of biological signaling pathways, enhancing its ability to learn and reason interactions within complex signal networks. We create prompts in the ProCoT (Protein Chain of Thought) format to incrementally decompose the signal transduction process like Definition 2, simulating real pathways of signal propagation. Specifically, we clarify the structure knowledge within the signaling pathway, such as signaling proteins and effector proteins. We then design rules for signal transmission across different levels to simulate the iterative process of signal transduction in proteins. Finally, we also design signal interruptions to simulate the continuity of protein transmission in real life. We use hard code to convert protein interaction information as Definition 1 in dataset into ProCoT format natural language to imitate protein signaling pathways as Definition 2. Figure 7 in the appendix is an example of ProCoT. The answer to this is "The relationship is activation".
@@ 3.1.2 Training on ProCot Format Dataset
In biology, co-expression and co-localization refer to the phenomenon in which proteins that are often expressed or located together in the cell tend to participate in the same or interconnected biological processes (Zhang et al., 2019). Thus, biologists frequently use known protein information to infer unknown protein interactions. Based on the principles of co-expression and co-localization in biology, we have formulated our training strategy. After constructing our ProCoT training data using DFS, we will obtain a circular protein interaction as in Figure 4. We mask the relationship between the initial and final proteins in the signaling pathway and let our model predict this relationship.
@@
3.1.3 Why ProCot Works: Biological Intuition behind ProCot
In this section, we will explore ProCoT through the lenses of biology and AI. (1) Simulating Biological Signaling Pathways: The biological signaling pathway is a series of ordered, rule-based interaction processes. During the process, the output of each step serves as the input for the next step, forming a highly organized information transmission chain. This pattern is consistent with the methods of handling serialized information in natural language processing (NLP). Additionally, Flan-T5’s self-attention blocks can capture the indirect relationships between proteins that are distant and do not have direct interactions. It is crucial for understanding biological signal transduction processes that involve multiple steps and complex intermediate links. (2) Signal Interruption Mechanism: This mechanism aims to mimic protein adaptability since the signaling between proteins needs to be constantly interrupted to ensure the cell’s accurate response to external changes. This mechanism we designed can satisfy the complex feedback mechanisms and regulatory networks within the cell. Overall, our ProCoT design follows biological principles.
@@ 3.2 Enhancing Protein Sequence and Function Comprehension
In our methodology, we enhance the model’s understanding of protein sequence by replacing the T5 embedding with the ProTrans embedding vectors and also perform instruction fine-tuning to enable the model to learn protein functions.